OpenStudy (anonymous):

Will FAN AND MEDAL!!!!!!!!!!!!!! Use elimination to find the solution to the system of equations.

OpenStudy (anonymous):

4x+y=10 3x-2y=2

OpenStudy (anonymous):

@triciaal Plz help this is due in a few i will fan and medal

OpenStudy (triciaal):

multiply eqn 1 by 2 then add new eqn with original eqn 2 you have one eqn with x solve then substitute in original to find y

OpenStudy (anonymous):

Theres 10 questions! :(

OpenStudy (triciaal):

so pay attention and do them the same process

OpenStudy (anonymous):

4x+y=10 3x-2y=2 heres the questions

OpenStudy (triciaal):

you also need to put some in new thread

OpenStudy (anonymous):

i had to ADDDDDDDDDDDDDDDDDD

OpenStudy (triciaal):